Solve the integral of x²cosxdx, using integration by parts.
1
Expert's answer
2021-06-09T16:37:10-0400

"\int x^2cos(x)dx=x^2sin(x)- \int sin(x).2xdx=x^2.sin(x)-2\int x.sin(x)dx"

"u=x^2 \space \space \space du=2x \space dx"

"dv=cos(x)dx \space \space \space v=sin(x)"

"x^2sin(x)-2(x.(-cos(x)- \int-cos(x)dx)=x^2sin(x)-2(-xcos(x)+ \int cos(x)dx"

"u=x \space \space \space \space du=dx"

"dv=sin(x)dx \space \space \space \space \space v=-cos(x)"

"=x^2.sin(x)+2xcos(x)-2sin(x)+C"